A stone is tied to a string (length = 0.552 m) and whirled in a circle at the same constant speed in two different ways. First, the circle is horizontal and the string is nearly parallel to the ground. Next, the circle is vertical. In the vertical case the maximum tension in the string is 5.30% larger than the tension that exists when the circle is horizontal. Determine the speed of the stone.
Answer #1

Given that,

Length = 0.552 m

When whirled Vertically,

T2 = m * g + m *v2 / r

When whirled Horizontally,

T1 = m * v2 / r

m * g + m * v2/ r = 1.053 * m * v2 / r

0.053 * mv2 / r = m * g

0.053 * v2 / r = g

v2 = 9.8 * 0.552 / 0.053

= 102.06

v = 1.43 m / s

Therefore, the speed of the stone = 1.43 m / s